OpenStudy (anonymous):

11. Without graphing, is the system independent, dependent, or inconsistent? y=-x+5 -3x-3y=-15 dependent independent inconsistent

OpenStudy (tyteen4a03):

Rebuild the first equation in the x y c form: x + y = 5 -3x -3y = -15 As you can see, the two equations are parallel (the second one is formed by multiplying the first equation by -3).

OpenStudy (anonymous):

so how do i tell if its dependent, independent, or inconsistent?????????????????

OpenStudy (anonymous):

@tyteen4a03

OpenStudy (tyteen4a03):

@sh9down45 I actually have no idea what exactly it's asking... let me google.

OpenStudy (anonymous):

ok, could u possibly help me with some other questions?

OpenStudy (anonymous):

these lines are parallel and overlap. so they are consistent because they have a solution, and dependent because there are infinitely many solutions

OpenStudy (tyteen4a03):

After some googling, it led me to this: http://www.purplemath.com/modules/systlin2.htm An independent system has exactly one intersection point (i.e one solution). A dependent system has infinite solutions (as the two lines overlap). An inconsistent system consists of parallel lines. They have the same slope but they do not overlap. This kind of system has no solutions. Since the second equation can be reduced to the first one, it is a dependent system.

OpenStudy (anonymous):

omg thank you so much can u help me with other questions? please!

OpenStudy (tyteen4a03):

@sh9down45 Close this current question and ask a new one. Please also choose @binarymimic as the best response, he responded quicker than me :P